OpenStudy (anonymous):

Write the equation of the line for the following condition. parallel to the line y = 4/9x + 5 through point (-2, 1) @shamil98

OpenStudy (shamil98):

parallel lines have the same slope but you gotta find b plug your values into y = mx + b and solve for b.

OpenStudy (anonymous):

So wait, what would be the point for y = 4/9x + 5

OpenStudy (shamil98):

no..

OpenStudy (shamil98):

(x,y) --> (-2,1) m = 4/9 plug them into y = mx + b and find the y-intercept.. (the value of b)

OpenStudy (anonymous):

1 = (4/9) - 2 + b correct?

OpenStudy (shamil98):

yes now solve for b

OpenStudy (anonymous):

17/9 ?

OpenStudy (shamil98):

yes

OpenStudy (shamil98):

so what is your final equation then?

OpenStudy (anonymous):

1 = (4/9) - 2 + 17/9

OpenStudy (shamil98):

no.. you want it in y =mx + b form you simply just put your slope and b value in y = (4/9)x + 17/9

OpenStudy (anonymous):

Oh wait so I don't add the (-2, 1) in the equation?

OpenStudy (shamil98):

nope.

OpenStudy (anonymous):

Oh, okay, I thought I did.